# Eliminating Backdoors in Neural Code Models via Trigger Inversion

### v1

@@ 5.3. Evaluation Results
Table 1 shows the performance of the baseline DBS and our EliBadCode in eliminating backdoors in 9 NCMs (= 3 model architectures * 3 tasks) used for the three code understanding tasks, i.e., defect detection, clone detection, and code search. Columns titled “Undefended” display the performance of the nine backdoored NCMs without any defense. Column titled “Average” presents the average values of the scores obtained by DBS and EliBadCode on CodeBERT, CodeT5, and UniXcoder. From this table, it is observed that, for the attack CodePoisoner, DBS has almost no effect in removing backdoors in nine NCMs. As described in Section 5.1, DBS cannot be applied to code search tasks. For defect detection tasks, EliBadCode can effectively reduce the average ASR from 99.04% to 0.93%, with almost no impact on the model’s normal predictive accuracy (64.14% vs. 63.40%). For clone detection tasks, EliBadCode decreases the average ASR from 100% to 4.73% significantly, while even improving the average ACC from 94.51% to 96.70%. For code search tasks, EliBadCode can increase the average ANR from 9.55 to 24.95, while maintaining the same average ACC. The backdoor attacks in NCMs for code search tasks aim to improve the ranking of the code snippet with the trigger given a query containing the target word. It is important to note that an ANR of 9.55 indicates that the backdoored model can elevate a (potentially malicious) code snippet injected with a trigger from its original rank at the 50% position to the 9.55% position. Assuming there are 100 candidate code snippets, 9.55% means that the trigger-injected code snippet would be ranked in the 10th position. In existing code search techniques (Sun et al., 2023), it is common practice to return the top 10 retrieved code snippets. Therefore, code snippets ranked in the top 10 are likely to be adopted by developers. Once the malicious trigger-injected code snippet is adopted and integrated into their projects, it can pose serious security risks. Although EliBadCode does not increase the 9.55% back to the original 50% position, it significantly reduces the risk of developers adopting the malicious trigger-injected code snippet.
For the attack BadCode, DBS has a certain backdoor removal effect, but the purified NCMs still exhibit a relatively high ASR. For example, for the defect detection task, the average ASR of the model purified by DBS still reaches 46.38%. This suggests that single-token triggers are easier to inversion compared to multi-token triggers, but they are more challenging to unlearn completely from the model. Compared to DBS, EliBadCode can significantly decrease the average ASR for defect detection and clone detection tasks to 2.64% and 6.11%, respectively. In addition, EliBadCode demonstrates strong backdoor elimination performance in code search tasks. Both DBS and EliBadCode maintain the model’s normal predictive performance (i.e., ACC) well while removing the backdoor. It means that removing backdoors by model unlearning effectively can contribute to maintaining the performance of models on clean data.
@@
For both attacks, it can also be observed that in clone detection tasks, the F1 scores of NCMs after removing backdoors are even higher than those of the backdoored NCMs without any defense. This is because fine-tuning NCMs with 10% of the trigger-injected training dataset does not negatively impact NCMs’ normal prediction performance; rather, the increased training data and training process enhance the NCMs’ effectiveness.
RQ2: Contribution of key designs in EliBadCode.
Table 2. Ablation study.
Method ACC ASR Epoch LD BLEU
w/o phase (a) 63.73% 100% - 9 12.36
w/o phase (b) 62.57% 0.24% 52 6 29.56
w/o Trigger Anchoring 60.98% 0.08% 25 15 26.33
EliBadCode 62.57% 0.24% 25 6 29.56
@@
To demonstrate the effectiveness of EliBadCode key design choices, we investigate the contributions of its various components. Table 2 presents the performance of EliBadCode on CodeBERT under the CodePoisoner attack with different designs. Rows 2–4 represent the performance of EliBadCode without phase (a): PL-specific trigger vocabulary generation, phase (b): sample-specific trigger position identification, and trigger anchoring in phase (c), respectively. Observe that without phase (a), the optimization search space expands, making it unable to invert the trigger close to the factual trigger. Therefore, the ASR results after model unlearning remain at 100%, unable to defend against backdoor attacks. Without phase (b) does not affect ASR or ACC of EliBadCode. As mentioned in Section 4.3, the purpose of the phase (b) design is to reduce the impact of adversarial perturbations and improve the efficiency of trigger inversion. Figure 10 shows the change of loss during the trigger optimization process for the defect detection task without phase (a) and phase (b). Observe that EliBadCode can invert a trigger close to the factual one by the 25th epoch, whereas without phase (b), it can only invert by the 52th epoch. EliBadCode with phase (b) is more efficient than that without phase (b) by two times. Without trigger anchoring, although the ASR of EliBadCode decreases, there is also a drop in ACC. As mentioned in Section 4.4, the purpose of the trigger anchoring design is to reduce the impact of noise tokens affect the prediction of the unlearned model on inputs containing them. Figure 10 shows the ACC of all test samples and the test samples containing noise tokens on undfended, w/o trigger anchoring and EliBadCode, respectively. Observed that EliBadCode achieves ACC close to the undefended results on both test samples. In contrast, w/o trigger anchoring achieves ACC of 60.98% and 54.92% on the two test samples, respectively, which are significantly lower than the undefended results. This indicates that the trigger anchoring in EliBadCode can effectively reduce the interference of noise tokens.
@@
To investigate the accuracy of trigger inversion, we also utilize three metrics to evaluate the difference between the inverted trigger and the factual trigger: Levenshtein Distance (LD) and BLEU (Papineni et al., 2002). LD represents the minimum number of edit operations required to transform one string into another. BLEU, a variant of the precision metric, calculates similarity by computing the n-gram precision of the inverted trigger compared to the factual trigger. The lower the LD and the higher the BLEU, indicate a higher precision of the inverted trigger. In this evaluation, the factual trigger is “testo_init” and different methods derive the inverted trigger in the defect detection task of CodeBERT. Table 3 shows the performance of the triggers inverted by baselines and EliBadCode. Observe that DBS has a very high LD (19/14) and very low BLEU (9.27/8.20). This indicates that the trigger inverted by DBS is significantly different from the factual trigger. EliBadCode achieves high precise in the inverted trigger, with an LD of only 6/5 and BLEU reaching 29.56/36.79, surpassing DBS by a significant margin. Additionally, in terms of LD and BLEU, EliBadCode outperforms w/o Trigger Anchoring (16/14, 19.62/23.30). This indicates that the inverted trigger with anchoring is closer to the factual trigger.
Table 3. Comparisons of inverted trigger and factual trigger. TA: Trigger Anchoring; DD: Defect Detection; CD: Clone Detection.
Task DBS w/o TG EliBadCode
@@
DD 19 9.27 16 19.62 6 29.56
CD 14 8.20 14 23.30 5 36.79
Figure 11. Influence of the numbers of clean samples.
RQ3: Influence of important settings, e.g., $k$ and $r$.
We study the influence of important settings on EliBadCode, including the number of clean samples in the trigger inversion, top-k and repeat size. It can be observed that the more clean samples there are, the better the performance of EliBadCode. In other words, the more clean samples there are, the more precise the trigger inverted by EliBadCode is, and the fewer epochs are needed. When the number is less than 20, EliBadCode cannot invert the correct trigger and thus cannot eliminate the backdoor in the model. When the number is 30, EliBadCode achieves the best results. Unfortunately, our experimental server can only support the input of up to 30 clean samples.
Top-k and repeat size are key parameters for EliBadCode in generating candidates during the GCG-based trigger inversion. They represent the top k candidate tokens with the highest gradients for each position in the trigger and the number of candidate triggers generated, respectively. Figure 13 and Figure 13 illustrate the impact of top-k and repeat size on the effectiveness of EliBadCode, respectively. It is worth noting that we fixed the top-k or repeat size at 64 to explore the effects of varying the other parameter on the effectiveness of EliBadCode. A smaller top-k will result in the factual trigger token not appearing among the candidate replacement tokens, while a larger top-k will reduce the probability of selecting the factual trigger token for replacement. A smaller repeat size will also reduce the probability of selecting the factual trigger token for replacement, while a larger repeat size will increase the time consumption of the trigger inversion. It can be observed that when both top-k and repeat size are 64, EliBadCode achieves the best performance with minimal time consumption.
Figure 12. Effect of number of candidate substitutes $k$.
Figure 13. Effect of times of repeat $r$.
RQ4: Performance of EliBadCode against adaptive attacks.
Table 4. Effectiveness of EliBadCode on adaptive attack.
@@
We study a scenario where the attacker understands the EliBadCode mechanism and attempts to bypass it. We design an adaptive attack targeting the GCG-based trigger inversion phase of EliBadCode. The idea is to encourage the injected trigger length (number of tokens) to be greater than the initialized trigger length (number of tokens) set by EliBadCode. Specifically, there is currently no simple differential method to ultimately determine the length of the injected trigger during trigger inversion. Therefore, we set the initialized trigger length to 5, which can cover more than 90% of identifier lengths (as described in Section 5.1). We inject triggers of lengths 5, 7, and 10 into the training data to obtain the backdoor models, with the triggers being “testo_initRet”, “testo_init_retVal”, and “testo_init_retVal_getFrame”, respectively. Other parameter settings are the same as in the RQ1 settings. According to Table 4, EliBadCode remains effective against renaming backdoor attacks with triggers longer than the set length. This is because EliBadCode can reverse engineer the effective part of the injected trigger, which can still use to effectively eliminate the backdoor in the model through trigger unlearning. It can be observed that as the injected trigger length increases, the defense effectiveness of EliBadCode gradually decreases. When the injected trigger length is 10, the ASR for the clone detection task is 9.16%, which may allow an attacker to launch a successful backdoor attack. However, as shown in Figure 5, identifiers with 10 tokens are very rare, and such long trigger data can be easily recognized as abnormal by developers (Sun et al., 2023). Therefore, it is difficult for attackers to bypass EliBadCode by increasing the length of the injected trigger.
6. Conclustion

### v2

@@ 5.3.1. RQ1: Effectiveness of EliBadCode in eliminating backdoors
Effectiveness of EliBadCode in eliminating backdoors. Table 1 shows the performance of the three baselines and our EliBadCode in eliminating backdoors in 27 NCMs. Columns titled “Undefended” display the performance of the 27 backdoored NCMs without any defense. Observe that for the attack CodePoisoner, on the defect detection and clone detection tasks, after applying ONION, the ASR remains high, ranging from 76.98% to 90.00% depending on the different models and tasks. On the code search task, ONION can increase the ANR to 18.25-21.56 and is better than AttDef (16.57-18.75). DBS has almost no effect in removing backdoors in 27 NCMs. Although AttDef significantly reduces the ASR in certain attack scenarios, it also greatly compromises the model’s normal performance. For example, on the defect detection task and CodeT5 architecture, AttDef can reduce the ASR to 18.64%, but it also lowers the ACC to 56.37%. It can also be observed that existing defenses excel at defending against different types of attacks. For example, AttDef and DBS perform comparably on token-level trigger-based attacks such as BadCode, outperforming ONION overall. However, ONION performs better than DBS on the identifier-level trigger-based attack CodePoisoner. Compared with the three baselines, on the defect detection task, our EliBadCode can significantly reduce the ASR to 0.15%-2.39%, depending on the different model architectures, while maintaining the ACC without a noticeable decline. Similarly, on the code clone task, EliBadCode can significantly decrease the ASR to 3.16%-5.86% while maintaining a high F1. On the code search task, EliBadCode can increase the ANR from 9.11-10.04 to 24.76-25.12, outperforming ONION and AttDef while maintaining the same average MRR. The backdoor attacks in NCMs for code search tasks aim to improve the ranking of the code snippet with the trigger given a query containing the target word. It is important to note that an ANR of 9.11 indicates that the backdoored model can elevate a (potentially malicious) code snippet injected with a trigger from its original rank at the 50% position to the 9.11% position. Assuming there are 100 candidate code snippets, 9.11% means that the trigger-injected code snippet would be ranked in the 10th position. In existing code search techniques, it is common practice to return the top 10 retrieved code snippets (Sun et al., 2023). Therefore, code snippets ranked in the top 10 are likely to be adopted by developers. Once the malicious trigger-injected code snippet is adopted and integrated into their projects, it poses serious security risks. Although EliBadCode does not increase the 9.11% back to the original 50% position, it significantly reduces the risk of developers adopting the malicious trigger-injected code snippet.
@@
For attacks BadCode and AFRAIDOOR, in all attack scenarios, EliBadCode continues to demonstrate the same excellent backdoor elimination capabilities and ability to maintain normal performance as it does for attack CodePoisoner, outperforming the two baselines. For all three attacks, we can observe that on clone detection tasks, the F1 scores of NCMs after removing backdoors are even higher than those of the backdoored NCMs without any defense. This is because fine-tuning NCMs with 10% of the trigger-injected training dataset does not negatively impact NCMs’ normal performance; rather, the increased training data and process enhance the NCMs’ effectiveness.
We conduct additional statistical tests to evaluate the differences between EliBadCode and the best baseline DBS. Using Prism software (Software, 1995), we compare the ASR results of DBS and EliBadCode across different tasks, models, and backdoor attack scenarios (a total of 18 comparisons). For each comparison, we perform an unpaired Wilcoxon-Mann-Whitney test (Wilcoxon et al., 1963) on all ASR scores for DBS and EliBadCode at a significance level of 5%. The results show that all p-values are $<0.0001$, indicating that EliBadCode significantly outperforms DBS.
Table 2. Impact of the generic fine-tuning under the attack CodePoisoner.
Attack Task Metric CodeBERT CodeT5 UniXCoder
CodePoisoner Defect Detection ACC 63.54% 64.12% 65.34%
ASR 98.80% 92.83% 92.25%
Clone Detection F1 96.56% 96.24% 97.34%
ASR 98.17% 96.23% 96.52%
Code Search MRR 0.81 0.81 0.81
ANR 15.15 16.17 15.22
Table 3. Impact of the generic fine-tuning under the attack BadCode.
Attack Task Metric CodeBERT CodeT5 UniXCoder
BadCode Defect Detection ACC 63.10% 64.04% 65.17%
ASR 98.13% 96.23% 96.17%
Clone Detection F1 96.53% 96.30% 97.31%
ASR 98.24% 96.71% 96.43%
Code Search MRR 0.81 0.81 0.81
ANR 15.78 16.11 15.38
In addition, considering part of our approach involves a fine-tuning to backdoored NCMs to make them forget the mapping between triggers and target labels. We further compare the effect of such a fine-tuning against that of a “generic” fine-tuning involving the same amount of data/training time to highlight the contribution given by the model unlearning. Specifically, we fine-tune poisoned NCMs using clean samples (i.e., without inverted backdoor triggers) while adhering to the same settings as the model unlearning (i.e., using 10% of the training data and identical hyperparameters). The experimental results shown in Table 3 and Table 3 demonstrate that the generic fine-tuning with clean data alone is ineffective in mitigating backdoors in backdoored NCMs. For instance, in the Defect Detection task, the ASR of the fine-tuned NCMs remains above 90%, indicating that backdoor behavior persists.
Table 4. Performance on clean models. Column “Clean” shows the performance of the original clean model.
Task Metric CodeBERT CodeT5 UniXCoder
Clean DBS EliBadCode Clean DBS EliBadCode Clean DBS EliBadCode
Defect Detection ACC 62.41% 62.19% 62.30% 64.17% 63.01% 63.18% 65.56% 64.27% 64.67%
Clone Detection F1 93.34% 95.65% 96.17% 94.67% 96.43% 96.18% 95.14% 97.29% 97.23%
Code Search MRR 0.81 0.81 0.81 0.81 0.82 0.82
Performance of EliBadCode on the clean NCMs. As mentioned in Section 4.4, after iterating over all possible target labels and producing a set of loss values and associated inverted triggers, one for each label. EliBadCode runs the outlier detection to obtain the ground-truth target label and associated inverted trigger. If the outlier detection yields a result, it indicates that the model is backdoored; otherwise, it is clean. Therefore, it is necessary to ensure that when the input is a clean model, EliBadCode can identify it and does not negatively impact its normal performance.
To investigate the performance of EliBadCode on clean NCMs, we test it on 24 models (6 clean and 18 backdoored) to assess its ability to distinguish between clean and backdoored models. Our results show that the false positive rate of DBS is 50%, while our EliBadCode does not produce false positives and can effectively differentiate between clean and backdoored models. In addition, we utilize the inverted trigger on the clean NCM to perform trigger unlearning, and the normal performance of the NCM results are shown in Table 4. From this table, it is observed that compared with the performance of the original clean model, the normal performance of the model fine-tuned with the inverted trigger dataset does not show a significant drop. This indicates that regardless of whether the input NCM is clean or backdoored, even if EliBadCode’s outlier detection is inaccurate, the trigger unlearning does not negatively impact the model’s normal performance.
5.3.2. RQ2: Contribution of key designs in EliBadCode
@@
Table 5 presents the performance of EliBadCode on CodeBERT under the CodePoisoner attack with different designs. Rows 2–4 represent the performance of EliBadCode without phase (a): PL-specific trigger vocabulary generation, phase (b): sample-specific trigger position identification, and trigger anchoring in phase (c), respectively. Observe that without phase (a), the optimization search space expands, making it unable to invert the trigger close to the factual trigger. Therefore, the ASR results after unlearning remain at 100%, unable to eliminate the backdoor. Without phase (b) does not affect ASR or ACC of EliBadCode. As mentioned in Section 4.3, the purpose of phase (b) is to reduce the impact of non-backdoor perturbations and improve the efficiency of trigger inversion. Figure 10 shows the change of loss during the trigger optimization process for the defect detection task without phase (a) and phase (b). Observe that EliBadCode can invert a trigger close to the factual one by the 25th epoch, whereas without phase (b), it can only invert by the 52nd epoch. EliBadCode with phase (b) is more efficient than that without phase (b) by two times. Without trigger anchoring, although the ASR of EliBadCode decreases, there is also a drop in ACC. As mentioned in Section 4.4, the trigger anchoring aims to reduce the impact of noise tokens on the prediction of the unlearned model on inputs containing them. Figure 10 shows the ACC of all test samples and the test samples containing noise tokens on undefended and EliBadCode without (w/o)/with trigger anchoring, respectively. Observed that EliBadCode with trigger anchoring achieves ACC close to the undefended results on both test sample sets. In contrast, w/o trigger anchoring achieves ACC of 60.98% and 54.92% on the two test sample sets, respectively, which are significantly lower than the undefended results. This indicates that the trigger anchoring in EliBadCode can effectively reduce the interference of noise tokens.
Table 5. Ablation study. TA: Trigger Anchoring.
Method ACC ASR
w/o phase (a) 63.73% 100%
w/o phase (b) 62.57% 0.24%
w/o TA 60.98% 0.08%
EliBadCode 62.57% 0.24%
@@
[…]
Influence of trigger anchoring.
To investigate the accuracy of trigger inversion, we also utilize two metrics to evaluate the difference between the inverted trigger and the factual trigger: Levenshtein Distance (LD) and BLEU (Papineni et al., 2002). LD represents the minimum number of edit operations required to transform one string into another. BLEU calculates similarity by computing the n-gram precision of the inverted trigger compared to the factual trigger. The lower the LD and the higher the BLEU, the higher the accuracy of the trigger inversion. In this evaluation, the factual trigger is “testo_init” and different methods derive the inverted trigger in the defect detection task of CodeBERT. Table 7 shows the performance of the triggers inverted by DBS and EliBadCode. Observe that DBS has a very high LD (19/14) and very low BLEU (9.27/8.20). This indicates that the trigger inverted by DBS is significantly different from the factual trigger. EliBadCode achieves high precision in the inverted trigger, with an LD of only 6/5 and BLEU reaching 29.56/36.79, surpassing DBS by a significant margin. Additionally, in terms of LD and BLEU, EliBadCode outperforms w/o Trigger Anchoring (16/14 and 19.62/23.30). This indicates that the inverted trigger with anchoring is closer to the factual trigger.
Table 6. Comparisons of the inverted trigger and factual trigger. Levenshtein Distance (LD).
Task DBS w/o TA EliBadCode
@@
Defect Detection 19 9.27 16 19.62 6 29.56
Clone Detection 14 8.20 14 23.30 5 36.79
Table 7. Performance on adaptive attack.
@@
5.3.3. RQ3: Influence of important settings, e.g., $\epsilon$, $k$ and $r$.
Influence of the number of clean samples. From Figure 14, it can be observed that the more clean samples are available, the better the performance of EliBadCode. It indicates that the more clean samples there are, the more precise the trigger inverted by EliBadCode is, and the fewer epochs are needed. When the number is less than 20, EliBadCode cannot invert the correct trigger and thus cannot eliminate the backdoor. When the number is 30, EliBadCode achieves the best results. Unfortunately, our experimental server can only support the input of up to 30 clean samples.
Influence of $\epsilon$, $k$ and $r$. $\epsilon$ is used to control the times of iterations in the trigger inversion process. $\epsilon$ is empirically determined. We experiment with several different values, including 50, 100, and 150. Our experimental results show that 1) when the model architecture is CodeT5, EliBadCode with $\epsilon=50$ is unable to invert the trigger in certain attack scenarios, such as when the attack is CodePoisoner and the task is Defect Detection; 2) with $\epsilon=100$ or $\epsilon=150$, EliBadCode can successfully invert the trigger in all 27 attack scenarios. Therefore, to balance effectiveness and efficiency, we set $\epsilon$ to 100 uniformly. $k$ and $r$ are key parameters for EliBadCode in generating candidates during the GCG-based trigger inversion. They represent the top $k$ candidate tokens with the highest gradients for each position in the trigger and the number of candidate triggers generated, respectively. Figure 14 and Figure 14 illustrate the impact of $k$ and $r$ on the effectiveness of EliBadCode, respectively. It is worth noting that we fixed the $k$ or $r$ at 64 to explore the effects of varying the other parameters on the effectiveness of EliBadCode. A smaller $k$ will result in the factual trigger token not appearing among the candidate replacement tokens, while a larger $k$ will reduce the probability of selecting the factual trigger token for replacement. A smaller $r$ will also reduce the probability of selecting the factual trigger token for replacement, while a larger $r$ will increase the time consumption of the trigger inversion. It can be observed that when both $k$ and $r$ are 64, EliBadCode achieves the best performance with minimal time consumption.
Figure 11. Effect of numbers of clean samples.
[…]
Effect of numbers of clean samples.
Figure 12. Effect of candidate substitutes $k$.
[…]
Effect of candidate substitutes $k$.
Figure 13. Effect of the times of repeat $r$.
[…]
Effect of the times of repeat $r$.
Figure 14. Distribution of identifier lengths.
[…]
Distribution of identifier lengths.
5.3.4. RQ4: Performance of EliBadCode against adaptive attacks
@@
We study a scenario where the attacker understands the EliBadCode mechanism and attempts to bypass it. We design an adaptive attack targeting the GCG-based trigger inversion phase of EliBadCode. The idea is to encourage the injected trigger length (i.e., number of tokens) to be greater than the initialized trigger length set by EliBadCode. Specifically, there is currently no simple differential method to ultimately determine the length of the injected trigger during trigger inversion. Therefore, we set the initialized trigger length to 5, which can cover more than 90% of identifier lengths (as described in Section 5.1). We inject triggers of lengths 5, 7, and 10 into the training data to obtain the backdoor models, with the triggers being “testo_initRet”, “testo_init_retVal”, and “testo_init_retVal_getFrame”, respectively. Other parameter settings are the same as in the RQ1 settings. The experimental results in Table 7 demonstrate that EliBadCode remains effective against renaming backdoor attacks with triggers longer than the set length. This is because EliBadCode can invert the effective part of the injected trigger, which can still be used to effectively eliminate the backdoor in the model through trigger unlearning. It can also be observed that as the injected trigger length increases, the defense effectiveness of EliBadCode gradually decreases. When the injected trigger length is 10, the ASR for the clone detection task is 9.16%, which may allow an attacker to launch a successful backdoor attack. However, as shown in Figure 14, identifiers with 10 tokens are very rare, and such long trigger data can be easily recognized as abnormal by developers (Sun et al., 2023). Therefore, it is difficult for attackers to bypass EliBadCode by increasing the length of the injected trigger.
6. Discussion
6.1. Performance of EliBadCode on Code LLMs
Existing backdoor attacks against NCMs have not yet been validated for effectiveness on code LLMs. To evaluate the effectiveness of EliBadCode on code LLMs, we first perform a backdoor attack on a popular code LLM called StarCoder using the attack CodePoisoner, and then apply baselines and EliBadCode to detect the backdoor in the backdoored StarCoder. The specific version of StarCoder we use is StarCoderBase-1B. The performance of EliBadCode and baselines are shown in Table 9. It is observed that 1) CodePoisoner achieves significant success in attacking StarCoder; 2) EliBadCode can effectively eliminate backdoors from the backdoored StarCoder and outperforms three baselines. This indicates that EliBadCode has the capability to ensure the security of large NCMs, i.e., code LLMs.
Of course, applying the defense (including EliBadCode) to larger LLMs depends not only on its own capabilities but also on the availability of sufficient computational resources. Once sufficient resources are available, EliBadCode can employ parameter-efficient fine-tuning methods (e.g., LoRA (Hu et al., 2022)) to perform trigger unlearning.
Table 8. Performance of EliBadCode on code LLM. Undef.: Undefended.
Attack Task Metric StarCoder
Undef. ONION DBS AttDef EliBadCode
CodePoisoner DD ACC 61.68% 59.98% 60.71% 58.81% 60.65%
ASR 96.57% 53.05% 20.23% 28.34% 2.87%
CD F1 71.91% 69.73% 70.15% 65.43% 70.18%
ASR 100% 60.43% 30.13% 45.07% 5.17%
CS MRR 0.70 0.65 - 0.69 0.69
ANR 10.23 18.43 - 19.07 26.36
Table 9. Time overhead of EliBadCode. TI: Trigger Inversion; TU: Trigger Unlearning. Eli.: EliBadCode.
Attack Task Phase CodeBERT CodeT5 UniXCoder
DBS Eli. DBS Eli. DBS Eli.
CodePoisoner DD TI 4m58s 22m40s 8m19s 26m13s 5m03s 24m31s
TU 0m28s 0m25s 1m02s 1m04s 0m26s 0m24s
CD TI 9m02s 28m31s 8m14s 53m09s 9m35s 35m07s
TU 3m13s 3m14s 8m17s 8m10s 3m10s 3m14s
CS TI - 32m25s - 26m10s - 50m50s
TU - 10m58s - 30m18s - 10m57s
6.2. Time Overhead of EliBadCode
During the experiments, to better understand the time cost of EliBadCode, we also recorded the time spent on trigger inversion and trigger unlearning phases. The time records are reported in Table 9. The time overhead of the trigger inversion in EliBadCode ranges from about 23 to 53 minutes, depending on the specific task and model. Although this time cost is significantly higher than DBS, it results in a substantial improvement in backdoor elimination effectiveness. Moreover, trigger inversion is a one-time, offline task, and this time cost is acceptable when compared to the substantial time overhead involved in training models. The low cost of trigger inversion in EliBadCode can be attributed to the design of phase (a) and phase (b). In the future, we will further optimize trigger inversion to enhance efficiency.
We also discover that our EliBadCode can be combined with ONION and AttDef, which defend by identifying and removing suspicious (trigger) tokens with high perplexity or attribution scores, to further reduce ASR. EliBadCode’s inverted trigger tokens can improve their identification accuracy by comparing the vector similarity between user input tokens and inverted trigger tokens. We experiment with this combined approach on two attacks by setting a vector similarity threshold of 0.9, meaning input tokens with a similarity greater than 0.9 are filtered. The results show that this approach further reduces the ASR to 0%. Notably, in this approach, EliBadCode does not need to perform unlearning, which can further reduce EliBadCode’s runtime.
6.3. Potential Limitations of Our Work
In addition to the limitation of EliBadCode in the efficiency of trigger inversion mentioned in the previous section, we discuss other potential limitations of our work in this section.
Firstly, as illustrated in Section 3, our defense mainly focuses on using third-party trained models. In particular, similar to existing baseline methods, we assume that defenders have a few local clean samples. Accordingly, our method is not feasible without clean samples. Besides, we need to train a model for the scenarios using third-party datasets before conducting trigger inversion and follow-up defenses, which is computation- and time-consuming. We will further explore how to conduct trigger inversion under few/zero-shot settings in our future works.
Secondly, similar to trigger inversion-based defenses in NLP (Liu et al., 2022; Shen et al., 2022), the trigger inversion process in our method also relies on a white-box setting. Accordingly, it does not apply to black-box scenarios in which the defenders can only access the final output of the backdoored model. We also note that in practical applications, it is often feasible to derive a white-box surrogate model from a black-box model using distillation techniques, as demonstrated in existing research (Su and McMillan, 2024; Yang et al., 2024a; Sun et al., 2024b). Once a white-box surrogate is obtained, EliBadCode can be applied to mitigate backdoor vulnerabilities. We will continue the exploration of designing black-box trigger inversion in our future works.
7. Conclusion